# Towards shutdownable agents via stochastic choice

### v2

@@ Appendix D Proof
For all policies $\pi$ and meta-episodes $E$ consisting of multiple mini-episodes, if $\pi$ maximizes expected return in $E$ given our DREST reward function, then $\pi$ is maximally USEFUL and maximally NEUTRAL.
Here is a proof sketch. Because $10<\lambda<$ 1, the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor is always positive, so expected return across the meta-episode $E$ is strictly increasing in the expected fraction of available coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize this latter quantity, and hence are maximally USEFUL. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is the total ($\gamma$-discounted) value of coins collected divided by the maximum total ($\gamma$-discounted) value of coins collected conditional on the agent’s chosen trajectory-length. The agent’s observations do not allow it to distinguish between different mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And since the discount factor $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is strictly decreasing in $N_{e_{i}}(L=l)$ – the number of times the relevant trajectory-length has previously been chosen in the meta-episode – the agent maximizes expected overall return by equalizing the probabilities with which it chooses each available trajectory-length. Therefore, optimal policies are maximally NEUTRAL.
@@
$\mathbb{E}_{\pi,E}(R)=\sum_{i=1}^{n}\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$
@@
$\textrm{USEFULNESS}(\pi)=\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$
@@
(Equalizing probabilities increases expected return) For any maximally USEFUL policies $\pi$ and $\pi^{\prime}$, any meta-episode $E$ consisting of multiple mini-episodes, and any trajectory-lengths $x$ and $y$, if:
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\\ +\sum_{i=1}^{s-1}\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1-i)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-i)\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (1)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (2)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=y)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}$ (3)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x\lor L=y)=\\ Pr_{\pi,e_{s}}\{L=x\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}\bigg{)}\\ +Pr_{\pi,e_{s}}\{L=y\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}\bigg{)}$ (4)
@@
$\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (5)
@@
$\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (6)
@@
$\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\\ >\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (7)
@@
$\Delta\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}-Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\})\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}+Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}>0$ (8)
@@
$\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}+Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}\geq 0$ (9)

### v3

@@ Appendix D Proof that DReST-optimal policies are maximally USEFUL and maximally NEUTRAL
For all policies $\pi$ and meta-episodes $E$ consisting of more than one mini-episode, if $\pi$ maximizes expected return in $E$ given our DReST reward function, then $\pi$ is maximally USEFUL and maximally NEUTRAL.
Here is a proof sketch. Because $10<\lambda<$ 1, the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor is always positive, so expected return across the meta-episode $E$ is strictly increasing in the expected fraction of available coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize this latter quantity, and hence are maximally USEFUL. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is defined as the total ($\gamma$-discounted) value of coins collected divided by the maximum total ($\gamma$-discounted) value of coins collected conditional on the agent’s chosen trajectory-length. The agent’s observations do not allow it to distinguish between different mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And since the discount factor $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is strictly decreasing in $N_{e_{i}}(L=l)$ – the number of times the relevant trajectory-length has previously been chosen in the meta-episode – the agent maximizes expected overall return by equalizing the probabilities with which it chooses each available trajectory-length. Therefore, optimal policies are maximally NEUTRAL.
@@
$\mathbb{E}_{\pi,E}(R)=\\ \sum_{i=1}^{n}\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$
@@
$\textrm{USEFULNESS}(\pi)=\\ \sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$
@@
(Equalizing probabilities increases expected return) For any maximally USEFUL policies $\pi$ and $\pi^{\prime}$, any meta-episode $E$ consisting of more than one mini-episode, and any trajectory-lengths $x$ and $y$, if:
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\\ \mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\\ +\sum_{i=1}^{s-1}\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1-i)\\ -\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-i)\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (1)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (2)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=y)=\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}$ (3)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x\lor L=y)=\\ Pr_{\pi,e_{s}}\{L=x\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}\bigg{)}\\ +Pr_{\pi,e_{s}}\{L=y\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}\bigg{)}$ (4)
@@
$\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)=\\ \mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (5)
@@
$\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)=\\ \mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (6)
@@
$\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\\ \sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\\ >\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\$
$+\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (7)
@@
$\Delta\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\\ -Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\})\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\\ +Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}>0$ (8)
@@
$\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\\ +Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}\geq 0$ (9)